# METok: Multi-Stage Event-based Token Compression for Efficient Long Video Understanding

### v1

@@ Abstract
Recent advances in Video Large Language Models (VLLMs) have significantly enhanced their ability to understand video content. Nonetheless, processing long videos remains challenging due to high computational demands and the redundancy present in the visual data. In this work, we propose METok, a training-free, Multi-stage Event-based Token compression framework designed to accelerate VLLMs’ inference while preserving accuracy. METok progressively eliminates redundant visual tokens across three critical stages: (1) event-aware compression during vision encoding, (2) hierarchical token pruning in the prefilling stage based on semantic alignment and event importance, and (3) a decoding-stage KV Cache optimization that further reduces memory consumption. Our experiments on diverse video benchmarks demonstrate that METok achieves an optimal trade-off between efficiency and accuracy by dynamically selecting informative visual tokens. For instance, equipping LongVA-7B with METok realizes an 80.6% FLOPs reduction and 93.5% KV Cache memory savings, all while maintaining comparable or even superior accuracy. The code is available here.
@@
[…]
@@ 1 Introduction
Video Large Language Models(VLLMs) Cheng et al. (2024); Wang et al. (2024); Lin et al. (2023); Li et al. (2024c); Zhang et al. (2024c) have recently achieved remarkable success in various tasks such as video question-answering, temporal reasoning, and grounding. However, extending these models to efficiently understand long videos remains a major challenge. Current VLLMs typically encode each frame into hundreds of tokens, leading to prohibitive computational and memory overhead as sequence lengths scale. Beyond the resource burden, this visual token proliferation also introduces semantic redundancy, where repetitive or low-salience tokens obscure truly informative visual cues. As the number of frames increases, critical visual information becomes diluted, ultimately degrading both model efficiency and performance.
@@
Building upon these observations, we propose METok, a Multi-stage Event-based Token compression framework that integrates visual-text alignment with hierarchical token reduction, without requiring additional training. METok operates in three stages: During the vision encoding stage, it segments videos into temporally coherent events and identifies key visual tokens while heavily compressing non-key ones. In the prefilling stage, it performs layer-wise hierarchical token pruning guided by attention and event importance. Finally, METok discards visual tokens from the KV Cache starting at the pruning boundary used in prefilling. We validate METok across diverse models and benchmarks, demonstrating that it can reduce computational cost by up to 80.6% and KV cache memory by over 90%, all while maintaining or even improving accuracy, and requiring no additional training. These three stages jointly improve memory and computational efficiency without compromising model performance, making METok an effective solution for long video understanding in VLLMs. The main contributions are summarized as follows:
@@ 2 Related Work
2.1 Video Large Language Models
The rapid growth of Large Language Models (LLMs)  Bai et al. (2023); Chiang et al. (2023); Touvron et al. (2023a) has further accelerated the development of VLLMs by providing a strong language backbone that can be extended with vision encoders. To extend these capabilities from images to videos, recent works have proposed various strategies for integrating temporal visual inputs into LLMs. LLaVA-OneVisionLi et al. (2024a) unifies image and video tasks within a single model that supports cross-modal transfer. LongVAZhang et al. (2024b) scales video comprehension by extrapolating LLM’s context length, allowing inference over thousands of visual tokens. However, most VLLMs still process video frames independently, encoding each frame into its own set of tokens. While suitable for short clips, this approach becomes prohibitively expensive for long videos due to the quadratic complexity of self-attention.
@@ 2.2 Token Compression
Token compression techniques Bolya et al. (2022); Zhong et al. (2024); Fu et al. (2024c); Yang et al. (2024); Wan et al. (2024); Liu et al. (2024); Shen et al. (2024); Tao et al. (2024) in VLLMs can be categorized by the inference stage at which they operate. Methods such as VisionZip Yang et al. (2024) and DivPrune Alvar et al. (2025) compress visual tokens early in the vision encoding stage. Prefilling-stage token compression approaches like FastV Chen et al. (2024) remove redundant visual tokens after text interaction begins, typically by analyzing attention weights within LLM to drop those contributing the least to multimodal reasoning. While these techniques Zhong et al. (2024); Tu et al. (2024); Xiao et al. (2023) bring efficiency gains, they often operate in isolation, overlooking the cumulative impact of redundancy across stages. In contrast, METok integrates token compression across all stages with spatiotemporal awareness, enabling efficient and effective processing of long video sequences.
@@ 3.1 Preliminary
Decoding stage. The decoding process follows an autoregressive pattern, in which each token is predicted sequentially by referencing previously generated tokens and stored key-value (KV) pairs.
@@ 3.2 METok
METok follows a structured three-stage approach tailored for long-video inference to progressively remove redundant visual tokens, achieving an optimal balance between computational efficiency and preservation of spatiotemporal semantics throughout the entire pipeline, as shown in Figure 2.
@@ 3.2.1 Vision Encoding: Event-Aware Token Reduction
As shown in Figure 3, a lower similarity score indicates a significant change in content, suggesting a potential event boundary. We select the $k{-}1$ lowest $S_{i}$ scores as event boundaries, dividing the video into $k$ events. Let $E$ denote the set of events.
@@
We average these scores per event and rank all $k$ events accordingly. The top $\lceil\alpha\cdot k\rceil$ events ($10<\alpha<$ 1) are designated as key events $E_{\text{key}}$ while the remaining are considered non-key events $E_{\text{non-key}}$. Within each event, whether key or non-key, METok further selects the top $\beta$ proportion of frames ($10<\beta<$ 1) that convey richer textual semantics as key frames, treating all remaining frames as non-key frames.
@@ 3.2.2 Prefilling: Hierarchical Token Pruning
where $r$ is a pruning factor ($10<r<$ 1), and $\alpha$ is the proportion of key events.
@@ 4 Experiments
Model #Frames FLOPs MVBench Egoschema MLVU VideoMME (w/o sub.) VideoMME (w sub.) (TB) Overall Long Overall Long Duration 16 sec 179.8 sec 3$\sim$120 min 1$\sim$60 min 30$\sim$60 min 1$\sim$60 min 30$\sim$60 min LLaVA-OneVision-0.5B (Li et al., 2024a) 32 3.9 45.5 26.8 50.3 44.0 35.4 43.5 38.7 FastV Chen et al. (2024) 32 2.2 45.0 26.5 50.8 42.4 35.8 46.4 39.4 VisionZip (Yang et al., 2024) 32 1.7 41.4 27.3 44.4 37.6 35.6 41.9 36.1 DivPrune Alvar et al. (2025) 32 1.2 44.2 26.1 37.4 42.6 36.0 46.4 39.5 METok (Ours) 32 1.3 44.7 25.8 52.1 42.7 36.4 47.0 39.7 LLaVA-OneVision-7B (Li et al., 2024a) 32 71.4 56.7 60.1 64.7 58.2 48.5 61.2 51.4 FastV Chen et al. (2024) 32 39.0 57.2 60.1 64.0 58.2 48.4 61.8 51.6 VisionZip (Yang et al., 2024) 32 27.7 49.4 56.4 55.0 49.6 44.6 55.2 48.4 DivPrune Alvar et al. (2025) 32 20.9 55.9 58.2 60.8 54.4 46.6 58.9 50.8 METok (Ours) 32 19.8 57.3 59.8 64.7 58.4 49.8 61.7 52.8 LongVA-7B (Zhang et al., 2024b) 128 241.6 50.8 43.5 59.0 52.6 46.2 54.3 47.6 FastV Chen et al. (2024) 128 133.6 50.9 44.2 59.2 52.9 45.6 55.7 47.0 VisionZip (Yang et al., 2024) 128 107.9 46.3 36.2 52.0 45.4 37.9 51.4 41.0 DivPrune Alvar et al. (2025) 128 50.1 50.6 43.9 58.7 52.4 45.9 55.8 47.4 METok (Ours) 128 46.8 50.7 44.4 60.4 52.4 46.6 56.0 47.7 VILA-1.5-13B (Lin et al., 2024) 16 84.1 50.9 50.4 50.4 49.5 42.4 53.3 46.7 FastV Chen et al. (2024) 16 49.8 50.4 50.0 49.5 48.4 41.6 53.1 45.9 VisionZip (Yang et al., 2024) 16 29.1 45.8 47.6 47.1 43.8 39.4 49.4 44.9 DivPrune Alvar et al. (2025) 16 20.3 50.7 50.4 50.3 48.9 41.8 53.5 46.3 METok (Ours) 16 19.6 50.5 50.5 50.5 49.5 42.7 53.9 48.1
@@ 4.1 Implementation Details
We evaluate METok using various open-source VLLMs spanning small to large scales, including LLaVA-OneVision-(0.5B, 7B) Li et al. (2024a), LongVA-7B Zhang et al. (2024b) and VILA-1.5-13B Lin et al. (2024), respectively sampling $32$, $128$, and $16$ frames. LLaVA-OneVision and VILA-1.5 adopt a SigLIP-pretrained ViT-L as their vision tower, while LongVA employs a CLIP-pretrained ViT-L. To assess computational efficiency, we measure the total FLOPs of both the prefilling and decoding stages using calflops Ye (2023). Prefill Time refers to the latency required to generate the first token. More implementation details are included in the Appendix C.
@@ 4.2 Main Results
Comparison with base models. Table 1 presents the performance of METok on LLaVA-OneVision-(0.5B, 7B), LongVA-7B and VILA-1.5-13B across representative benchmarks. METok consistently achieves substantial reductions in FLOPs while maintaining or improving accuracy, demonstrating a trade-off between efficiency and performance. In particular, METok shows clear advantages in long-context scenarios. On the long subset of VideoMME, it consistently surpasses all baselines, indicating its ability to retain essential temporal and semantic visual tokens under extended durations. The results also highlight METok’s scalability across model sizes, from LLaVA-OneVision-0.5B to VILA-1.5-13B, achieving up to 80.6% FLOPs reduction without degrading performance.
@@ 4.3 Ablation Study
Temporal Event Segmentation. We also compare various event segmentation strategies on LLaVA-OneVision-7B. As shown in Figure 5, uniform segmentation lowers FLOPs to 27.9% of the baseline but noticeably hurts performance, while random segmentation further reduces FLOPs to 21.1% yet also degrades accuracy. In contrast, our temporal segmentation strategy strikes the best balance by reducing FLOPs to 27.7% while outperforming the other methods on both MLVU and VideoMME.
@@ 5 Conclusion
In this paper, we introduced METok, a training-free multi-stage token compression framework that significantly enhances VLM efficiency for long-video understanding. By integrating event-aware segmentation, hierarchical token pruning, and KV Cache optimization, METok substantially reduces computational costs while preserving accuracy. Our extensive experiments on LLaVA-OneVision-(0.5B, 7B), LongVA-7B and VILA-1.5-13B show that METok achieves up to 80.6% FLOPs reduction and 93.5% KV cache savings with comparable performance. METok generalizes well to different model sizes and architectures, while also improving memory consumption for the tested VLLMs.
@@
While recent progress has introduced video large language models (VLLMs) exceeding 13B parameters, our current evaluation is limited to models up to 13B (e.g., VILA-1.5-13B), due to computational resource constraints. These larger models often require substantial GPU memory and infrastructure, which are beyond our reach at this stage. Nonetheless, we conduct comprehensive ablation studies across multiple open-source VLLMs and benchmarks, and the consistent improvements validate the general effectiveness of METok. We leave the evaluation on larger models to future work as resources become available.
@@ Acknowledgements
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b211dd. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Additionally, this work also benefited from the scientific support and HPC resources provided by The Hessian Center for Artificial Intelligence (hessian.AI).
@@ Appendix A Datasets Details
  • Video-MME Fu et al. (2024a): Video-MME is a large-scale benchmark designed to comprehensively assess the video understanding capability of MLLMs. It contains 900 videos spanning 254 hours in total, with durations ranging from 11 seconds to over 60 minutes. The dataset covers 6 high-level categories and 30 subcategories, and includes manually annotated multiple-choice questions that require spatial-temporal reasoning and multi-event analysis. Our evaluation follows the official test set in a zero-shot manner.
@@ Appendix B Pseudocode
2 $S_{i}\leftarrow\cos(v_{i},v_{i+1}),\quad\forall i\in[1,T{-}1]$;
3 $E\leftarrow\text{Split}(V)$ at $(k{-}1)$ lowest $S_{i}$ positions;
4
@@ Appendix C Implementation Details
For both LLaVA-OneVision-(0.5B, 7B) and LongVA-7B, we use $(s_{1},s_{2})=(2,3)$, $L_{H}=[3,10,19]$, and $\alpha=0.5$ as shared hyperparameters. The remaining parameters vary by model: for LLaVA-OneVision-(0.5B, 7B), we set $k=5$, $\beta=0.4$, and $r=0.76$; for LongVA-7B, we set $k=13$, $\beta=0.45$, and $r=0.55$. For VILA-1.5-13B, we use the setting of $(s_{1},s_{2})=(2,3)$, $L_{H}=[13,24,34]$, $\alpha=0.5$, $k=3$, $\beta=0.4$, and $r=0.65$.

### v2

@@ Abstract
Recent advances in Vision Large Language Models (VLLMs) have significantly enhanced their ability to understand video content. Nonetheless, processing long videos remains challenging due to high computational demands and the redundancy present in the visual data. In this work, we propose METok, a training-free, Multi-stage Event-based Token compression framework designed to accelerate VLLMs’ inference while preserving accuracy. METok progressively eliminates redundant visual tokens across three critical stages: (1) event-aware compression during vision encoding, (2) hierarchical token pruning in the prefilling stage based on semantic alignment and event importance, and (3) a decoding-stage KV Cache optimization that further reduces memory consumption. Our experiments on diverse video benchmarks demonstrate that METok achieves an optimal trade-off between efficiency and accuracy by dynamically selecting informative visual tokens. For instance, equipping LongVA-7B with METok realizes an 80.6% FLOPs reduction and 93.5% KV Cache memory savings, all while maintaining comparable or even superior accuracy. The code is available here.
@@
Mengyue Wang1,7   Shuo Chen2,6,7   Kristian Kersting3,4,5 Volker Tresp2,7   Yunpu Ma2,7 22footnotemark: 2 1Technical University of Munich   2 LMU Munich   3DFKI SAINT   4Hessian AI 5TU Darmstadt   6Konrad Zuse School of Excellence in Reliable AI (relAI) 7Munich Center for Machine Learning (MCML)
@@ 1 Introduction
Vision Large Language Models (VLLMs) Cheng et al. (2024); Wang et al. (2024); Lin et al. (2023); Li et al. (2024d); Liao et al. (2024a); Bi et al. (2025b) have recently achieved remarkable success in various tasks such as video question-answering, temporal reasoning, and grounding. However, extending these models to efficiently understand long videos remains a major challenge. Current VLLMs typically encode each frame into hundreds of tokens, leading to prohibitive computational and memory overhead as sequence lengths scale. Beyond the resource burden, this visual token proliferation also introduces semantic redundancy, where repetitive or low-salience tokens obscure truly informative visual cues. As the number of frames increases, critical visual information becomes diluted, ultimately degrading both model efficiency and performance.
@@
Building upon these observations, we propose METok, a Multi‐stage Event‐based Token compression framework that integrates visual‐text alignment with hierarchical token reduction, without requiring additional training. METok operates in three stages: During the vision encoding stage, it segments videos into temporally coherent events and identifies key visual tokens while heavily compressing non-key ones. In the prefilling stage, it performs layer-wise hierarchical token pruning guided by attention and event importance. Finally, METok discards visual tokens from the KV Cache starting at the pruning boundary used in prefilling. We validate METok across diverse models and benchmarks, demonstrating that it can reduce computational cost by up to 80.6% and KV cache memory by over 90%, all while maintaining or even improving accuracy, and requiring no additional training. These three stages jointly improve memory and computational efficiency without compromising model performance, making METok an effective solution for long video understanding in VLLMs. The main contributions are summarized as follows:
@@ 2 Related Work
2.1 Vision Large Language Models
The rapid growth of Large Language Models (LLMs)  Bai et al. (2023); Chiang et al. (2023); Touvron et al. (2023a) has further accelerated the development across a wide range of areas Liao et al. (2024b); Zhang et al. (2025, 2024b); Bi et al. (2025a) and notably in VLLMs, where strong language backbones are extended with vision encoders. To extend these capabilities from images to videos, recent works have proposed various strategies for integrating temporal visual inputs into LLMs. LLaVA-OneVisionLi et al. (2024a) unifies image and video tasks within a single model that supports cross-modal transfer. LongVAZhang et al. (2024b) scales video comprehension by extrapolating LLM’s context length, allowing inference over thousands of visual tokens. However, most VLLMs still process video frames independently, encoding each frame into its own set of tokens. While suitable for short clips, this approach becomes prohibitively expensive for long videos due to the quadratic complexity of self-attention.
@@ 2.2 Token Compression
Token compression techniques Bolya et al. (2022); Zhong et al. (2024); Fu et al. (2024c); Yang et al. (2024); Wan et al. (2024); Liu et al. (2024); Shen et al. (2024); Tao et al. (2024); Xing et al. (2025) in VLLMs can be categorized by the inference stage at which they operate. For instance, AuroraCap Chai et al. (2025) merges similar visual tokens within transformer layers, LVC Wang et al. (2025) introduces a parameter-free query-attention video compression mechanism, St3 Zhuang et al. (2025) prunes inattentive visual tokens progressively across LLM layers, VideoChat-Flash Li et al. (2024c) segments video and applies hierarchical compression in two stages, and FastVid Shen et al. (2025) adopts a density-based token pruning strategy to maintain essential information. Methods such as VisionZip Yang et al. (2024) and DivPrune Alvar et al. (2025) compress visual tokens early in the vision encoding stage. Prefilling-stage token compression approaches like FastV Chen et al. (2024) remove redundant visual tokens after text interaction begins, typically by analyzing attention weights within LLM to drop those contributing the least to multimodal reasoning. While these techniques Zhong et al. (2024); Tu et al. (2024); Xiao et al. (2023) bring efficiency gains, they often operate in isolation, overlooking the cumulative impact of redundancy across stages. In contrast, METok integrates token compression across all stages with spatiotemporal awareness, enabling efficient and effective processing of long video sequences.
@@ 3.1 Preliminary
Decoding stage. The decoding process follows an autoregressive pattern, in which each token is predicted sequentially by referencing previously generated tokens and stored key‐value (KV) pairs.
@@ 3.2 METok
METok follows a structured three‐stage approach tailored for long‐video inference to progressively remove redundant visual tokens, achieving an optimal balance between computational efficiency and preservation of spatiotemporal semantics throughout the entire pipeline, as shown in Figure 2.
@@ 3.2.1 Vision Encoding: Event-Aware Token Reduction
As shown in Figure 3, a lower similarity score indicates a significant change in content, suggesting a potential event boundary. We select the k1k{-}1 lowest $S_{i}$ scores as event boundaries, dividing the video into $k$ events. Let $E$ denote the set of events.
@@
We average these scores per event and rank all $k$ events accordingly. The top $\lceil\alpha\cdot k\rceil$ events ($0<\alpha<1$) are designated as key events $E_{\text{key}}$ while the remaining are considered non-key events $E_{\text{non-key}}$. Within each event, whether key or non‐key, METok further selects the top $\beta$ proportion of frames ($0<\beta<1$) that convey richer textual semantics as key frames, treating all remaining frames as non‐key frames.
@@ 3.2.2 Prefilling: Hierarchical Token Pruning
where $r$ is a pruning factor ($0<r<1$), and $\alpha$ is the proportion of key events.
@@ 4 Experiments
[…]
@@ 4.1 Implementation Details
We evaluate METok using various open-source VLLMs spanning small to large scales, including LLaVA-OneVision-(0.5B, 7B) Li et al. (2024a), LongVA-7B Zhang et al. (2024b) and VILA-1.5-13B Lin et al. (2024), respectively sampling $32$, 128128, and $16$ frames. LLaVA-OneVision and VILA-1.5 adopt a SigLIP-pretrained ViT-L as their vision tower, while LongVA employs a CLIP-pretrained ViT-L. To assess computational efficiency, we measure the total FLOPs of both the prefilling and decoding stages using calflops Ye (2023). Prefill Time refers to the latency required to generate the first token. More implementation details are included in the Appendix C.
@@ 4.2 Main Results
Comparison with base models. Table 1 presents the performance of METok on LLaVA-OneVision-(0.5B, 7B), LongVA-7B and VILA-1.5-13B across representative benchmarks. METok consistently achieves substantial reductions in FLOPs while maintaining or improving accuracy, demonstrating a trade-off between efficiency and performance. In particular, METok shows clear advantages in long-context scenarios. On the long subset of VideoMME, it consistently surpasses all baselines, indicating its ability to retain essential temporal and semantic visual tokens under extended durations. The results also highlight METok’s scalability across model sizes, from LLaVA-OneVision-0.5B to VILA-1.5-13B, achieving up to 80.6% FLOPs reduction without degrading performance. Notably, for each base model, we use a single set of hyperparameters across all benchmarks without task-specific tuning. METok consistently delivers strong performance under this unified setting, demonstrating robustness to hyperparameter variation.
@@ 4.3 Ablation Study
Temporal Event Segmentation. We also compare various event segmentation strategies on LLaVA-OneVision-7B. As shown in Figure 5, uniform segmentation lowers FLOPs to 27.9% of the baseline but noticeably hurts performance, while random segmentation further reduces FLOPs to 21.1% yet also degrades accuracy. In contrast, our temporal segmentation strategy strikes the optimal balance by reducing FLOPs to 27.7% while outperforming the other methods on both MLVU and VideoMME benchmarks.
@@ 5 Conclusion
In this paper, we introduced METok, a training-free multi-stage token compression framework that significantly enhances VLLM efficiency for long-video understanding. By integrating event-aware segmentation, hierarchical token pruning, and KV Cache optimization, METok substantially reduces computational costs while preserving accuracy. Our extensive experiments on LLaVA-OneVision-(0.5B, 7B), LongVA-7B and VILA-1.5-13B show that METok achieves up to 80.6% FLOPs reduction and 93.5% KV cache savings with comparable performance. METok generalizes well to different model sizes and architectures, while also improving memory consumption for the tested VLLMs. Overall, METok offers a unified and efficient solution that focuses on maintaining long-range temporal and semantic coherence, making it particularly well-suited for long-form video tasks.
@@
While recent progress has introduced vision large language models (VLLMs) exceeding 13B parameters, our current evaluation is limited to models up to 13B (e.g., VILA-1.5-13B), due to computational resource constraints. These larger models often require substantial GPU memory and infrastructure, which are beyond our reach at this stage. Nonetheless, we conduct comprehensive ablation studies across multiple open-source VLLMs and benchmarks, and the consistent improvements validate the general effectiveness of METok. We leave the evaluation on larger models to future work as resources become available.
@@ Acknowledgements
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b273dd. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Additionally, this work also benefited from the scientific support and HPC resources provided by The Hessian Center for Artificial Intelligence (hessian.AI), as well as the the Federal Ministry for Research, Technology and Space (BMFTR) project "XEI: Extremely Efficient Inference for Large Context Length" (XEI), project identification number 01IS24079B. This paper is also supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Research, Technology and Space.
@@ Appendix A Datasets Details
  • Video-MME Fu et al. (2024a): Video-MME is a large-scale benchmark designed to comprehensively assess the video understanding capability of VLLMs. It contains 900 videos spanning 254 hours in total, with durations ranging from 11 seconds to over 60 minutes. The dataset covers 6 high-level categories and 30 subcategories, and includes manually annotated multiple-choice questions that require spatial-temporal reasoning and multi-event analysis. Our evaluation follows the official test set in a zero-shot manner.
@@ Appendix B Pseudocode
2
32pt $S_{i}\leftarrow\cos(v_{i},v_{i+1}),\quad\forall i\in[1,T{-}1]$;
4 $E\leftarrow\text{Split}(V)$ at (k1)(k{-}1) lowest $S_{i}$ positions;
5
@@ Appendix C Implementation Details
We use a single set of hyperparameters for each base model across all benchmarks without task-specific tuning. For both LLaVA-OneVision-(0.5B, 7B) and LongVA-7B, we use $(s_{1},s_{2})=(2,3)$, $L_{H}=[3,10,19]$, and $\alpha=0.5$ as shared hyperparameters. The remaining parameters vary by model: for LLaVA-OneVision-(0.5B, 7B), we set $k=5$, $\beta=0.4$, and $r=0.76$; for LongVA-7B, we set $k=13$, $\beta=0.45$, and $r=0.55$. For VILA-1.5-13B, we use the setting of $(s_{1},s_{2})=(2,3)$, $L_{H}=[13,24,34]$, $\alpha=0.5$, $k=3$, $\beta=0.4$, and $r=0.65$.